OpenStudy (anonymous):

The area of the parallelogram shown is thirty-six millimeters squared. What is the perimeter?

Parth (parthkohli):

Area of a parallelogram is base * height. A possibility here is 4 and 9. Perimeter would then be 4 + 4 + 9 + 9 = 8 + 18 = 26